OpenStudy (anonymous):

Both James and Fred brought in a can of chicken noodle soup for lunch. James' can of soup has a diameter of 5 inches and a height of 6 inches. Fred's can of soup has a diameter of 6 inches and a height of 5 inches. Who has the bigger can of soup - Fred or James?

OpenStudy (anonymous):

volume of the soup = pi * r^2* H(eight) so james' soup = pi*5*(6/2)^2 = pi*5*9 = 45*pi fred's soup = pi*6*(5/2)^2 = 75/2*pi = 37.5*pi james' can is bigger.